QUESTION 3: [25]

A [a] m x [b] m double-pane window consists of two [c]-mm-thick layers of glass 

(k = [d] W/m·K) that are separated by a [e] mm air gap (kair = [f] W/m·K). The heat flow 

through the air gap is assumed to be by conduction. The inside and outside air 

temperatures are [g]°C and [h]°C, respectively, and the inside and outside heat 

transfer coefficients are [i] and [j] W/m2

·K. 

3.1) Draw a representation of the setup with the thermal resistance network (5)

3.2) Determine the daily rate of heat loss through the window in steady operation. (15)

3.3) Calculate the temperature difference across the largest thermal resistance. (


1
Expert's answer
2021-05-17T00:53:01-0400



"\dfrac{dQ}{dT} = \dfrac{kA∆T}{l} = \dfrac{A∆T}{\frac lk}"


"\dfrac{dQ}{dT} = \dfrac{A∆T}{\frac{l_1}{k_1} +\frac{l_2}{k_2}+ \frac{l_3}{k_3} }"